# Eye-gaze Guided Multi-modal Alignment Framework for Radiology

### v2

@@ IV Experiments
In this study, we conduct zero-shot classification on multiple datasets, as well as image-to-text and text-to-image retrieval tests. In the following subsections, we first introduce the implementation details in Section IV-A. Then, we compare our proposed framework with state-of-the-art medical vision-language pre-training methods and show the comparison results in Section IV-B. Further, we analyze different components in our framework through ablation study in Section IV-C. Finally, the feature representation is visualized in Section IV-D.
IV-A Implementation Details
We employed the MIMIC-EYE [14] dataset as our training set, consisting of 3689 images extracted from the MIMIC datasets [30, 5, 31, 32]. Each sample is accompanied by corresponding eye-tracking data and transcripts text. These eye-tracking data are provided by the publicly available EYE GAZE [33] and REFLACX [34] datasets on PhysioNet [35]. Additionally, we evaluate the model’s zero-shot classification and zero-shot retrieval capabilities using the CheXpert [6], RSNA [36], and SIIM-ACR [37] datasets individually. Detailed descriptions of the datasets can be found in the supplementary materials. We use SwinTransformer [38] as the image encoder, BioClinicalBERT [39] as the text encoder, and initialized the model parameters using pre-training parameters provided by the MedCLIP model, which were trained on CheXpert [6] and MIMIC-CXR [5] datasets. It’s worth noting that the two CheXpert subsets used in our test set are not included in the aforementioned training set, thus avoiding any data leakage issues. Subsequently, we continued training the model on MIMIC-EYE. For the transcribed text in the data, we first split it into individual sentences. After passing through the text encoder, we concatenated the features of sentences belonging to the same report and computed the similarity with image features. As images need to correspond to the positions of gaze attention maps, we only employed contrast-related image augmentation methods during training. And we train our model with 50 epochs with an initial learning rate $1\times 10^{-6}$ and weight decay $1\times 10^{-4}$ and 10 epochs of warm-up. All eye-gaze heatmaps and images are simply resized to $224\times 224$ pixels.
IV-B Comparison with State-of-the-Arts
IV-B1 Zero-shot Classification
In this section, we compare the diagnostic performance and generalization of our EGMA model with other prominent models across three datasets CheXpert 5x200 [6], RSNA [36], and SIIM-ACR [37] for zero-shot classification. These models include the universal image-text contrastive model CLIP [1] and three SOTA multi-modal methods in the medical domain: GLoRIA [10], MedCLIP [9], and MGCA [11]. In our testing, the textual queries we used are all derived from [10], written by a board-certified radiologist. We evaluate the common metrics of accuracy (Acc.) and F1 score (F1) for the classification task.
@@ IV-B1 Zero-shot Classification
TABLE I: Comparison results of zero-shot classification tasks with other SOTA models on CheXpert 5x200, RSNA, and SIIM-ACR datasets. Bold denotes the best result and Underline denotes the second-best result.
@@
MGCA [11] 43.60 41.37 60.83 57.77 30.03 25.45
EGMA(Ours) 60.50 58.44 57.81 36.70 63.62 61.46
The experimental results are presented in Tab. I, demonstrating competitive performance of our model. For the CLIP model in the first row of Tab. I, its training data mainly consist of natural images, hence its processing capability for medical images is relatively poor. Conversely, models in rows two to four utilize encoders pre-trained on medical datasets, thus, they are expected to outperform CLIP, as evident in the results on the CheXpert 5x200 and RSNA datasets. And our model achieves the best performance on the CheXpert 5x200 and SIIM-ACR datasets, and the second-best performance on the RSNA dataset. This could be attributed to significant differences between the RSNA dataset and CheXpert 5x200/SIIM-ACR in terms of pathological features related to pneumonia, image characteristics, and annotation methods. Additionally, the RSNA dataset may suffer from class imbalance issues, which adversely affect model performance. MGCA exhibits the best performance in pneumonia diagnosis on the RSNA dataset. This could be attributed to its large training dataset, MIMIC-CXR, which contains a significant number of pneumonia-related cases. However, as indicated in the visualization analysis in Sec. IV-D, the feature representation of MGCA is inferior to that of our model. Interestingly, both the GLoRIA and MGCA models show lower diagnostic performance for pneumothorax disease on the SIIM-ACR dataset compared to the CLIP model. This could be attributed to the influence of the training data distribution during the pre-training of these models, leading to the learning of irrelevant shortcut features [12] in the dataset and resulting in over-fitting and decreased generalization ability. Consequently, when testing with datasets that differ from the pre-training data, such as the SIIM-ACR dataset, performance degradation is observed, as shown in Tab. I. In contrast, our model does not experience this phenomenon in these datasets. This is attributed to the introducing of eye-gaze data, which contains prior information from radiologists. Guided by this information, our model can learn the universal features of texts and images truly relevant to the disease in the data, enabling effective discrimination between normal and diseased samples even when applied to different datasets. Specifically, our model achieves the best performance on the CheXpert 5x200 and SIIM-ACR datasets, and the second-best performance on the RSNA dataset.
IV-B2 Image-text Retrieval
TABLE II: Comparison results of zero-shot retrieval task with other SOTA models on CheXpert 8x200 dataset. Bold denotes the best result and Underline denotes the second-best result. We do not report the Image-to-text retrieval results of ConVIRT [18] because it has not released code yet.
@@ IV-B2 Image-text Retrieval
To further validate the alignment capability of our model between texts and images, we compare the zero-shot retrieval performance of EGMA with other models on CheXpert 8x200 dataset [18]. Following [10], for each disease category in this dataset, we have five corresponding textual prompts. It is worth noting that the prompts for retrieval tasks are different from those for classification tasks in the previous section, but all are written by board-certified radiologists. In the image-to-text retrieval task, we first compute the similarity between the image and all candidate texts, and then rank the retrieved results. Similarly, in the text-to-image task, we compute the similarity between the textual prompts and all images, and rank the retrieval results. We report Precision at Top-1, Top-5, and Top-10, which reflect how many relevant examples are retrieved. As shown in Tab. II, our model achieves the best results in both retrieve tasks. Our model outperforms the second-best model in the image-to-text and text-to-image retrieval tasks by 3.9%, 5.88%, and 4.33%, and 19.75%, 14.50%, and 12% in terms of P@1, P@5, and P@10 metrics, respectively. This indicates that our model has fully learned the relationship between images and texts, achieving better alignment effects.
IV-C Ablation Study
@@
TABLE III: Comparison results of zero-shot classification ablation experiments on CheXpert 5x200, RSNA, and SIIM-ACR datasets. Bold denotes the best result and Underline denotes the second-best result.
Method CheXpert5x200 RSNA SIIM-ACR
@@
Unified(Ours) 60.50 58.48 57.81 36.70 63.62 61.46
1% Gaze 59.00 56.68 40.39 29.74 57.65 57.09
5% Gaze 59.00 56.76 52.94 35.13 59.33 58.62
10% Gaze 58.70 56.32 44.19 31.57 58.40 57.85
30% Gaze 59.70 58.80 51.46 34.61 58.76 58.46
50% Gaze 59.60 58.83 58.01 36.82 61.38 58.85
To further validate the model’s performance, we conducted ablation experiments on the proposed EGF and EGM modules, while also assessing the impact of the proportion of eye-gaze data on the model results. As shown in the upper half of Tab. III, the first row represents our Baseline model, where we utilize the initialized weights from the MedCLIP model. The second row “MLCE” indicates that within our EGF module, the EGF loss is not further computed beyond the Eq. 5, instead, only the multi-label cross-entropy (MLCE) loss between the eye-gaze guided similarity matrix and the model’s output similarity matrix is calculated. The third row “EGF” utilizes the Eye-gaze Guided Fine-grained loss described in Eq. 6. The fourth row “EGM” indicates that the model is trained solely through the Eye-gaze Guide cross-model Mapping method. Finally, the fifth row presents our proposed EGMA model, which integrates the aforementioned modules guided by eye-gaze data. In Tab. III, it can be observed that the method using only gaze-guided MLCE loss significantly improves performance compared to the baseline on CheXpert 5x200 dataset, with a slight improvement on RSNA but a severe decline on SIIM-ACR dataset. However, models using EGF or EGM show significant improvements on SIIM-ACR. This indicates that while MLCE improves performance on some datasets, it simultaneously reduces the model’s generalization ability. Thus, relying solely on simple loss for similarity matrix is insufficient. In this work, by combining eye-gaze guided image-text relationships with fine-grained feature alignment, although the model’s performance slightly decreases on CheXpert 5x200, its overall generalization improves. Similarly, to enhance the model’s multi-modal alignment ability, introducing eye-gaze guided cross-modal mapping results in improved performance and generalization, with EGM achieving optimal performance on RSNA dataset. Finally, when optimizing both fine-grained alignment and cross-modal alignment using eye-gaze, the model achieves optimal performance on all three datasets, demonstrating further enhancement in generalization.
Numerous studies [25, 28, 13, 40] have demonstrated that training models using eye-gaze data can achieve comparable performance to models trained with fine-grained manual annotations. Meanwhile, the cost of collecting fine-grained manual annotations is significantly higher than that of collecting eye-gaze data. Therefore, incorporating eye-gaze into pre-training tasks is a feasible approach to enhancing model performance. To further validate the efficiency of our model using eye-gaze data, we conduct ablation experiments on the proportion of it in the training set. Our training dataset, MIMIC-EYE, consists of a total of 3695 samples. We perform five experiments using 1%, 5%, 10%, 30%, and 50% of the eye-gaze data, resulting in 37, 185, 370, 1108, and 1848 samples with prior information from radiologists, respectively. As shown in the lower part of Tab. III, the model’s performance on the CheXpert 5x200 dataset improved when trained with 1% of eye-gaze data. However, due to the limited data volume, the model’s performance on other datasets is inferior to the baseline. When increasing the eye-gaze data to 5%, the model shows significant improvements on all three datasets. With the continuous increase in eye-gaze data, the performance of the model also improves. Therefore, even with a small amount of eye-gaze data, our framework can effectively guide the model’s multi-modal processing capability, ensuring performance enhancement. This further illustrates the reliability and applicability of our model.
IV-D Visualization of Feature Representation
V Limitation
The collection of eye-gaze data has long been a challenging task in the medical domain. Previous collection systems [42, 43, 28] suffered from limitations such as restricting head movement of radiologists and the inability to manipulate medical images (e.g., adjust contrast, scale size), resulting in low efficiency in data collection. However, recent efforts [44, 13] have introduced more flexible collection systems, allowing radiologists to manipulate images as they would in their routine work, minimizing the impact of data collection on them. Therefore, we believe that eye-gaze data collection can achieve a certain scale in the future and effectively alleviate the pressure of manual annotation. Furthermore, due to the limited availability of multi-modal data containing eye-gaze information, we plan to compare it with fine-grained annotated datasets from more sources in the future to further validate the auxiliary role of eye-gaze data.
VI Conclusion
@@
In this work, we reveal the significant role of radiologists’ eye-gaze data in multi-modal alignment and propose an Eye-gaze Guided Multi-modal Alignment framework called EGMA. Our framework first processes eye-gaze data into token-level relation matrices, then utilizes these matrices to optimize fine-grained alignment between image patches and text tokens. Furthermore, the framework integrates cross-modal mapping, leveraging eye-gaze data to guide feature mapping between images and texts bidirectionally, thereby enhancing the model’s ability to handle multi-modal data. We evaluate EGMA’s zero-shot capabilities on multiple datasets and observe significant performance improvements in classification and retrieval tasks. Additionally, we investigate the impact of eye-gaze data scale on performance, finding that even small amounts of eye-gaze data can enhance the model’s multi-modal alignment capabilities during pretraining. Our future work will continue to explore auxiliary information in eye-gaze data, such as temporal features, and further optimize their role in multi-modal feature alignment. We believe this work can serve as a valuable reference for the application of eye-gaze data in multi-modal frameworks and promote its development in the field of medical multi-modality.

### v3

@@ IV Experiments
In this study, we first conduct supervised and zero-shot classification as well as zero-shot retrieval experiments in Sec. IV-A to validate the model’s generalization performance and its representation capability of multi-modal features. Then, in Sec. IV-B, we perform ablation studies on various modules of EGMA. Additionally, to further investigate the auxiliary effect of eye-gaze data, we compare the performance when guided by different amounts of eye-gaze data. Finally, in Sec. IV-C, we visualize the model’s feature representations and the learned image-text relationships, further demonstrating the model’s performance and interpretability.
@@
TABLE I: Comparison results of supervised classification task with other SOTA models on CheXpert, RSNA, and SIIM-ACR datasets. Area under ROC curve (AUROC) is reported with different portions of training data: 1%, 10%, 100%. Red and blue denote the best and second-best results.
Method CheXpert[39] RSNA [40] SIIM-ACR [41]
1% 10% 100% 1% 10% 100% 1% 10% 100%
ConVIRT [24] 85.90 86.80 87.30 77.40 80.10 88.60 - - -
BioViL [25] 81.95 85.37 88.62 81.76 85.68 88.64 80.26 82.79 90.51
MedKLIP [26] - - - 87.31 87.99 89.31 85.27 90.71 91.88
MGCA [11] 85.80 87.66 89.30 85.22 87.54 89.24 86.12 89.66 92.16
GLoRIA [10] 86.60 87.80 88.10 86.10 88.00 88.60 - - -
PRIOR [43] 86.16 87.08 89.08 86.72 88.07 89.19 88.35 89.72 92.49
MedCLIP [9] 85.74 87.49 88.02 87.61 88.19 89.10 88.84 91.13 92.18
EGMA(Ours) 87.71 88.92 89.50 88.41 89.40 90.10 90.78 92.17 93.29
IV-A Comparison with State-of-the-Arts
Image Classification We conduct supervised classification experiments on the CheXpert [39], RSNA [40], and SIIM-ACR [41] datasets. CheXpert [39] is a large-scale public dataset for chest radiograph interpretation, it comprises 224,316 chest radiographic images. Following [11], we utilize the official training split as our training set, and the official validation set of 202 images with expert-label as our test set. RSNA [40] is a comprehensive dataset for Pneumonia diagnosing. It contains 29,700 chest X-ray images categorized into $normal$ and $pneumonia$ positive category. We follow [11] to divide the data into 70% for training, 15% for validation, and 15% for testing. SIIM-ACR [41] is a chest dataset used for pneumothorax diagnosing. It consists of 2379 images with pneumothorax and 8300 images without pneumothorax. In this work, we utilize a subset defined in [42] as our test set, with the remaining data used for training and validation. More details of dataset can be found in the supplementary materials.
In the supervised classification experiments, we adopt the linear classification settings [10], where the pre-trained image encoder is frozen, and only a randomly initialized linear classification head is trained. We adopt area under ROC curve (AUROC) metric to evaluate all model’s performance. And for better validate the model’s efficiency, we test its performance using 1%, 10%, and 100% of the training set. As shown in Tab. I, our model achieved the best results compared to other models. Additionally, with only 1% of the training set, our model outperformed the second-best model by 1.11%, 0.8%, and 1.94% on the CheXpert, RSNA, and SIIM-ACR datasets, respectively. Moreover, as the amount of training data increased, the model’s performance improved significantly. This demonstrates that, with the assistance of radiologists’ eye-gaze data, our model possesses strong multi-modal feature representation capabilities.
TABLE II: Comparison results of zero-shot classification tasks with other SOTA models on CheXpert 5x200, RSNA, and SIIM-ACR datasets. The Accuracy (Acc.) and F1-score (F1) metrics are reported. Red and blue denote the best and second-best results.
@@ IV-A Comparison with State-of-the-Arts
EGMA(Ours) 61.30 60.38 76.97 43.49 63.62 61.46
We further conduct zero-shot classification tasks on the CheXpert5x200 [10], RSNA [40], and SIIM-ACR [41] datasets. CheXpert5x200 includes five common chest diseases, $Atelectasis$, $Cardiomegaly$, $Consolidation$, $Edema$, and $Pleural\ Effusion$, each with 200 chest X-rays. It is important to note that the CheXpert training set does not include any data from CheXpert5x200, so there is no data leakage issue. The test sets for RSNA and SIIM-ACR are the same as those used in the supervised classification task. All text prompts are provided by a professional radiologist [10]. During testing, we calculated the similarity between image features and text prompt features for all diseases, with the highest similarity indicating the predicted category. As shown in Tab. II, CLIP [1] performs poorly on medical images due to its training data primarily consisting of natural images. The models in rows two to five use encoders pre-trained on medical datasets, and thus, their performance is better than that of CLIP. Interestingly, the GLoRIA [10] and MGCA [11] perform worse than the CLIP model in diagnosing pneumonia on the SIIM-ACR dataset. This indicates that these models are significantly influenced by the data distribution, resulting in poor generalization performance. Conversely, our EGMA achieves the best results in all other metrics, except for the F1-score on the RSNA dataset. This demonstrates that our model, enhanced by eye-gaze data, has learned more generalizable feature relationships between medical images and text, significantly improving its generalization performance.
TABLE III: Comparison results of zero-shot retrieval task with other SOTA models on CheXpert 8x200 dataset. The Precision at Top-1, Top-5, and Top-10 are reported. Red and blue denote the best and second-best results.
@@
Image-text Retrieval To further validate the alignment capability of our model between visual and textual features, we compare the zero-shot retrieval performance of EGMA with other models on CheXpert 8x200 dataset [24]. Unlike CheXpert5x200 [10], CheXpert8x200 includes eight common chest diseases, $No\ Finding$, $Cardiomegaly$, $Edema$, $Pneumonia$, $Atelectasis$, $Pneumothorax$, $Pleural\ Effusion$, and $Fracture$, each with 200 chest X-rays and five corresponding text prompts. It is worth noting that the prompts for retrieval tasks are different from those for classification tasks in the previous section, but all are written by board-certified radiologists. In the image-to-text retrieval task, we first compute the similarity between the image and all candidate texts, and then rank the retrieved results. Similarly, in the text-to-image task, we compute the similarity between the textual prompts and all images, and rank the retrieval results. We report Precision at Top-1, Top-5, and Top-10, which reflect how many relevant examples are retrieved. As shown in Tab. III, our model achieves the best results in both retrieve tasks. Our model outperforms the second-best model in the image-to-text and text-to-image retrieval tasks by 3.9%, 5.88%, and 4.33%, and 19.75%, 14.50%, and 12% in terms of P@1, P@5, and P@10 metrics, respectively. This indicates that our model has fully learned the relationship between images and texts, achieving better alignment effects.
IV-B Ablation Study
To further validate the model’s performance, we conducted ablation experiments on the proposed EGF and EGM modules, while also assessing the impact of the proportion of eye-gaze data on the model results. As shown in the upper half of Tab. IV, the first row represents our Baseline model, where we utilize the initialized weights pre-trained on CheXpert [39] and MIMIC-CXR [35] datasets [9]. The second row “MLCE” indicates that within our EGF module, the EGF loss is not further computed beyond the Eq. 4, instead, only the multi-label cross-entropy (MLCE) loss between the eye-gaze guided similarity matrix and the model’s output similarity matrix is calculated. The third row “EGF” utilizes the Eye-gaze Guided Fine-grained loss described in Eq. 5. The fourth row “EGM” indicates that the model is trained solely through the Eye-gaze Guide cross-model Mapping method. Finally, the fifth row presents our proposed EGMA model, which integrates the aforementioned modules guided by eye-gaze data.
In Tab. IV, it can be observed that the method using only gaze-guided MLCE loss significantly improves performance compared to the baseline on CheXpert 5x200 dataset, with a slight improvement on RSNA but a severe decline on SIIM-ACR dataset. However, models using EGF or EGM show significant improvements on SIIM-ACR. This indicates that while MLCE improves performance on some datasets, it simultaneously reduces the model’s generalization ability. Thus, relying solely on simple loss for similarity matrix is insufficient. In this work, by combining eye-gaze guided image-text relationships with fine-grained feature alignment (EGF), although the model’s performance slightly decreases on CheXpert 5x200, its overall generalization improves. Similarly, to enhance the model’s multi-modal alignment ability, introducing eye-gaze guided cross-modal mapping results in improved performance and generalization, with EGM achieving optimal performance on RSNA dataset. Finally, when optimizing both fine-grained alignment and cross-modal alignment using eye-gaze, the model achieves dominant performance on all three datasets, demonstrating further enhancement in generalization.
Numerous studies [16, 17, 14, 13] have demonstrated that training models using eye-gaze data can achieve comparable performance to models trained with fine-grained manual annotations. Meanwhile, the cost of collecting fine-grained manual annotations is significantly higher than that of collecting eye-gaze data. Therefore, incorporating eye-gaze into pre-training tasks is a feasible approach to enhancing model performance. To further validate the efficiency of our model using eye-gaze data, we conduct ablation experiments on the proportion of it in the training set. Our training dataset, MIMIC-EYE, consists of a total of 3695 samples. We perform ablation experiments using 1%, 5%, 10%, and 50% of the eye-gaze data, resulting in 37, 185, 370, and 1848 samples with prior information from radiologists, respectively. We repeat each experiment three times to eliminate the bias caused by random sampling, and report the average results. As shown in the lower part of Tab. IV, the model’s performance on the CheXpert 5x200 dataset improved when trained with 1% of eye-gaze data. However, due to the limited data volume, the model’s performance on other datasets is inferior to the baseline. When increasing the eye-gaze data to 5%, the model shows significant improvements on all three datasets. With the continuous increase in eye-gaze data, the performance of the model also improves. Therefore, even with a small amount of eye-gaze data (185 samples), our framework can effectively guide the model’s multi-modal processing capability, ensuring performance enhancement. This further illustrates the applicability of our model and its low training cost characteristics.
@@ IV-B Ablation Study
TABLE IV: Comparison results of zero-shot classification ablation experiments on CheXpert 5x200, RSNA, and SIIM-ACR datasets. The Accuracy (Acc.) and F1-score (F1) metrics are reported. Each value in the lower part is the average of three runs. Red and blue denote the best and second-best results.
Method CheXpert5x200 [39] RSNA [40] SIIM-ACR [41]
@@
Unified(Ours) 61.30 60.38 76.97 43.49 63.62 61.46
1% Gaze 58.93$\pm$0.06 56.62$\pm$0.05 40.38$\pm$0.01 29.75$\pm$0.01 57.90$\pm$0.21 57.37$\pm$0.24
5% Gaze 58.93$\pm$.006 56.69$\pm$0.06 53.00$\pm$0.05 35.13$\pm$0.01 59.20$\pm$0.11 58.51$\pm$0.10
10% Gaze 59.30$\pm$0.01 57.82$\pm$0.01 53.78$\pm$0.01 35.37$\pm$.001 58.27$\pm$0.11 57.71$\pm$0.12
50% Gaze 59.55$\pm$0.07 58.84$\pm$0.01 58.54$\pm$0.07 37.01$\pm$0.20 61.41$\pm$0.31 58.88$\pm$0.22
IV-C Visualization
Figure 3: Results of cross-modality attention maps visualization. Related text content: (a) ”heart size borderline enlarged”; (b) ”increased bibasilar opacities are the combination of increased bilateral pleural effusions and bibasilar atelectasis”.
Figure 4: t-SNE visualization on CheXpert 5x200 dataset by CLIP and our EGMA. The figures display points of different colors representing various ground truth disease types and their cluster assignments. The color-coded points illustrate the clustering results of each algorithm.
To better demonstrate the correspondence learned by the EGMA framework between text and radiographic images, we conducted a cross-modality attention maps visualization in Fig. 3. Guided by eye-gaze data, the EGMA framework clearly outperforms other state-of-the-art methods in the field in accurately localizing disease regions. In Fig. 4, we visualize the feature representations of CLIP [1] and our EGMA model on images of CheXpert 5x200 dataset using the t-SNE [44]. It can be observed that our model exhibits better clustering representation. The CLIP model, which was not trained on medical data, is unable to effectively differentiate these diseases. More results of t-SNE visualization can be referred to the supplementary materials, clustering performance of other SOTA methods [10, 11] also inferior to our EGMA.
V Discussion and Conclusion
@@
In this work, we reveal the significant role of radiologists’ eye-gaze data in multi-modal alignment and propose an Eye-gaze Guided Multi-modal Alignment framework called EGMA. Our framework first processes eye-gaze data into token-level relation matrices, then utilizes these matrices to optimize fine-grained alignment between image patches and text tokens. Furthermore, the framework integrates cross-modal mapping, leveraging eye-gaze data to guide feature mapping between images and texts bidirectionally, thereby enhancing the model’s ability to handle multi-modal data. We evaluate EGMA’s zero-shot capabilities and fine-tuned performances on multiple datasets and observe significant improvement in classification and retrieval tasks. Additionally, we investigate the impact of eye-gaze data scale on performance, finding that even small amounts of eye-gaze data can enhance the model’s multi-modal alignment capabilities during pre-training. Overall, our EGMA framework explores the feasibility of incorporating eye-gaze data from radiologists to assist in multi-modal feature alignment during model training, laying the foundation for the application of eye-gaze data in the medical multi-modal domain.
V-A Limitations and Discussion
Our work only compared state-of-the-art methods in classification and retrieval tasks, without conducting downstream tasks such as lesion localization or segmentation. Additionally, our model heavily relies on multi-modal datasets like MIMIC-EYE [33], which can simultaneously collect eye-gaze data, medical images, and diagnostic text. The scenarios for collecting these data are also a significant consideration. For instance, in clinical ultrasound diagnosis [32], radiologists often use both hands to operate the equipment and verbally communicate their diagnostic information to an assistant. In this context, it is convenient to simultaneously record ultrasound images, eye-gaze data, and audio. In contrast, during chest X-ray diagnosis in MIMIC-EYE, radiologists typically record diagnostic information directly in text form rather than verbally. Fortunately, some recent efforts [18, 14] are focusing on how to naturally collect multi-modal data of radiologists during diagnosing. They have designed more flexible collection systems that better accommodate the routine work of radiologists, which is crucial for the widespread adoption of collecting multi-modal diagnostic data such as eye-gaze information.
V-B Potential Impacts
Although the eye-gaze data we used is publicly available and we have permission to use it, some studies [45, 46] have indicated that private information such as gender, age, and mental state of observers can be extracted from eye-gaze data. Therefore, privacy concerns have always been a focal point in using eye-gaze data. To address this, we recommend using de-identification methods to filter eye-gaze data or releasing the data in the form of heatmaps rather than the raw data.
V-C Future Work
In the future, we will continue to optimize these proposed collection systems [45, 46] and explore the guidance role of eye-gaze data between images and handwritten diagnostic reports to accelerate their application in real medical diagnostic scenarios. This will provide a research foundation to alleviate data annotation pressure and enhance model interpretability. Additionally, we will continue to analysis the eye-gaze features, such as temporal features, and further optimize their role in multi-modal feature alignment. We believe this work can serve as a valuable reference for the application of eye-gaze data in multi-modal frameworks and promote its development in the field of medical multi-modality.